OpenStudy (ny,ny):

on a foraging trip, Grover dug up 7 bones and caught 4 fish in 49 minutes. Harold, working at the same speed as Grover, dug up 5 bones and caught 2 fish in 29 minutes. Under these conditions, how long does it take to dig up a bone, and how long does it take to catch a fish?

OpenStudy (anonymous):

First of all give a pro numeral value for fish and for bones. In this case x and y. The establish the two equations. 7x+4y=49 (1) 5x+2y=29 (2) Alter the second equation by multiplying it by two 10x+4y=58 (3) Now we can minus eqn (3) from (1) -3x=-9 x=3 Sub Back In: 7(3)+4y=49 21+4y=49 4y=28 y=7

OpenStudy (anonymous):

Had to edit a few times. I'm finished now. @NY,NY

OpenStudy (ny,ny):

okay, so how long does it take for each? because i can't say "it took 3 to dig up bones and 7 to catch a fish." what's the unit? hours?

OpenStudy (ny,ny):

it can't be hours.

OpenStudy (ny,ny):

@Henry.Lister what's the unit?

OpenStudy (ny,ny):

but... omg yur right thank yu =)

OpenStudy (ny,ny):

lol yes thanks Henry :D